OpenStudy (anonymous):

convert r=2sin3(theta) to rectangular form?? please:)

OpenStudy (anonymous):

haha i hav a test on all this stuff tom..so yeahh

OpenStudy (anonymous):

if z = r(cosθ + isinθ) z^3 = r^3 (cos3θ + isin3θ)

OpenStudy (anonymous):

or can we multiply r to both sides and get r^2=2rcos3(theta)?..n thn end up with x^2+y^2-6y=0?

OpenStudy (amistre64):

r = 2sin^3(t) right?

OpenStudy (anonymous):

how is sin 3θ equal to 6y?

OpenStudy (anonymous):

no\[r=2\sin3\Theta\]

OpenStudy (amistre64):

r = 2 sin(3t) then, ok

OpenStudy (anonymous):

i multiplied 2 times 3..u cant do tht huh

OpenStudy (anonymous):

no u cant

OpenStudy (anonymous):

oh..lol..k

OpenStudy (anonymous):

if z = r(cosθ + isinθ) = x + iy z^3 = r^3 (cos3θ + isin3θ) = (x^3 - 3xy^2) + i(3x^2 y - y^3)

OpenStudy (anonymous):

thats the eqn?

OpenStudy (anonymous):

no..long way to go

OpenStudy (anonymous):

equate the coefficients r^3sin3θ = (3x^2 y - y^3)

OpenStudy (anonymous):

r u indian?

OpenStudy (anonymous):

so sin3θ is (3x^2y - y^3)/r^3 now think back to ur original qstn we get r^4 = 2(3x^2y - y^3)

OpenStudy (anonymous):

m indian yes..

OpenStudy (anonymous):

wud r^4 b (x^2+y^2)^2

OpenStudy (anonymous):

yes maam

OpenStudy (anonymous):

wer in south india?

OpenStudy (anonymous):

i come from the extreme north but i live in the capital

OpenStudy (anonymous):

delhi

OpenStudy (anonymous):

yeah..nd m kashmiri

OpenStudy (anonymous):

oh..cool

OpenStudy (anonymous):

wht grade r u in?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

ohk...

OpenStudy (anonymous):

so uve done complex nos?

OpenStudy (anonymous):

not rlly..our teacher didint go over those tht well

OpenStudy (anonymous):

ohk...theyre really useful..did u get all those rotations i ws doing?

OpenStudy (anonymous):

yeah:)

OpenStudy (anonymous):

then u know complex quite well

OpenStudy (anonymous):

wht time is it there?

OpenStudy (anonymous):

haha..well..being asian helps with math

OpenStudy (anonymous):

its 4...der?

OpenStudy (anonymous):

2 in the morning

OpenStudy (anonymous):

oh wow..ohk..well i shud get back 2 all the othr probs:(

OpenStudy (anonymous):

:(..thanks for the medal..post ur probs..ill help:)

OpenStudy (anonymous):

no problem..thnkss

OpenStudy (anonymous):

ur welcome..keep in touch..